OpenStudy (anonymous):

Convert 5243 to base 2. I know I need to convert 5243 to base 10, then convert base 10 to base 2. But I don't know how to convert to base 10.

OpenStudy (amistre64):

what base is it in now?

OpenStudy (amistre64):

(427)_10 is what in base 2? well, divide by 2 and pull off the remainder and the remainders form the base 2 numeral starting at the right and working left either that or when you get to the end pull them off left to right 427 = 213(2) + 1 <--- right most digit 213 = 106(2) + 1 106 = 53(2) + 0 53 = 27(2) + 1 27 = 13(2) + 1 13 = 6(2) + 1 6 = 3(2) + 0 3 = 1(2) + 1 1 = 0(2) + 1 <--- left most digit (427)_10 = (110111011)_2